Presented below are financial statements (except cash flows) for two not-for-profit organizations. Neither organization has any permanently restricted net assets.

ABC Not-for-Profit XYZ Not-for-Profit
Statement of Activities Unrestricted Temporarily
Restricted
Unrestricted Temporarily
Restricted
Revenues
Program service revenue $ 5,685,000 $ 2,178,200
Contribution revenues 3,380,900 $ 745,500 3,097,900
Grant revenue 95,400 $ 765,300
Net gains on endowment investments 17,800
Net assets released from restriction
Satisfaction of program restrictions 457,800 (457,800 ) 364,900 (364,900 )
Total revenues 9,541,500 383,100 5,641,000 400,400
Expenses
Education program expenses 5,712,400 1,521,100
Research program expense 1,356,000 2,201,100
Total program service expenses 7,068,400 3,722,200
Fund-raising 469,700 365,000
Administration 730,000 1,235,000
Total supporting service expenses 1,199,700 1,600,000
Total expenses 8,268,100 5,322,200
Increase in net assets 1,273,400 383,100 318,800 400,400
Net assets January 1 4,308,000 735,000 774,700 197,700
Net assets December 31 $ 5,581,400 $ 1,118,100 $ 1,093,500 $ 598,100
STATEMENT OF NET ASSETS ABC Not-for-Profit XYZ Not-for-Profit
Current assets
Cash $ 65,400 $ 171,300
Short-term investments 165,500 62,400
Supplies inventories 20,400 94,500
Receivables 152,700 55,800
Total current assets 404,000 384,000
Noncurrent assets
Pledges receivable 625,000
Long-term investments 2,790,000
Land, buildings, and equipment (net) 2,140,100 2,058,000
Total noncurrent assets 5,555,100 2,058,000
Total assets $ 5,959,100 $ 2,442,000
Current liabilities
Accounts payable $ 23,000 $ 122,500
Total current liabilities 23,000 122,500
Noncurrent liabilities
Notes payable 178,500
Total noncurrent liabilities 178,500
Total liabilities 201,500 122,500
Net Assets
Unrestricted 5,715,000 1,391,400
Temporarily restricted 42,600 928,100
Permanently restricted 0 0
Total net assets 5,757,600 2,319,500
Total liabilities and net assets $ 5,959,100 $ 2,442,000

Required:
a. Calculate the following ratios (Assume 365 days):

  • Program expense.
  • Fund-raising efficiency.
  • Working capital.


b. For each ratio, which of the two organizations has the stronger ratio.
(Do not round intermediate calculations. Round "Program expense" answers to 1 decimal place and "Fund-raising efficiency" answers to 3 decimal places and "Working capital" answers to nearest whole number.)
Answer #1
Formula ABC XYZ
A) Program Expense Program Expense/ Total expense (=7068400/8268100) 0.8549 (=3722200/5322200) 0.699372

It determines how the expense related to a project are related to total expenses of the company.

B) Fund raising ratio determines how efficiently fund has been raised by the company. It is calculated by the following formula:

Fund raising ratio= Unrestricted contribution/ Fund raising expenses

ABC Co= 5,71,5000/ 469700 = 12.16

XYZ co = 1391400/365000= 3.81

C) Working Capital Ratio = It is difference of current assets and current liabilities.

ABC Co = 404000-23000 = 381000

XYZ co= 384000-122500 = 261500